A 200,000 kg railroad car moving at 5.14 m/s couples up (sticks together) with another 150,000 kg car at rest. What is the final speed of the two cars? [Write answer to 3 sig. figs.]
Answer #1

using conservation of momentum

m1 v1 + m2 v2 = ( m1 + m2) v

2* 5.14 + 0 = ( 1.5 + 2) v

v = 2.943 m/s

=====

Comment before rate in case any doubt, will reply for sure.. goodluck